Zorba Zantani runs a business, Zantani Manufacturers. Zantani Manufacturers, manufactures kitchen utensils and sells them to retailers who, in turn, sells the product to the public.

Zantani Manufacturers is a registered VAT vendor and only deals with other VAT vendors.

All amounts include VAT at 15%, where applicable, unless otherwise indicated.

In February 2019, Zantani Manufacturers manufactured 1 500 potato peelers at a cost of R690 each (including VAT).

The labour cost to produce the 1 500 potato peelers amounted to R25 000.

1. Calculate the cost price of manufacturing the potato peelers incurred by Zantani Manufacturers during February 2019. (6)

2. What type of account is VAT Input, an asset, a liability, an expense or an income? (1)

3. Calculate the amount of VAT Input claimable by Zantani Manufacturers during February 2019. (4)

Zantani Manufacturers has a mark-up of 30% on cost and sells all the potato peelers to Outside Retailers.




1
Expert's answer
2021-05-11T15:09:27-0400

Q1) cost price = cost of manufacturing 1500patato peelers + cost of labour

= (690 * 1500) + 25000

= 1035000 +25000

= 1060000


Q2) Liability Account


Q3) VAT = 15% of cost of production - labour

= 1060000-25000=1035000

= 1035000* 15/100

= 155250

Q4) Markup =30%

markup = selling price-cost of production

cost of production =70% = 1060000


70%=1060000

100%=?

=1060000*100/70

106000000/70= 1514285.7

selling price = 1514285.7

markup = 1514285.7- 1060000

= 454285.7